# TEOChat: A Large Vision-Language Assistant for Temporal Earth Observation Data

### v1

@@ Abstract
Large vision and language assistants have enabled new capabilities for interpreting natural images. These approaches have recently been adapted to earth observation data, but they are only able to handle single image inputs, limiting their use for many real-world tasks. In this work, we develop a new vision and language assistant called TEOChat that can engage in conversations about temporal sequences of earth observation data. To train TEOChat, we curate an instruction-following dataset composed of many single image and temporal tasks including building change and damage assessment, semantic change detection, and temporal scene classification. We show that TEOChat can perform a wide variety of spatial and temporal reasoning tasks, substantially outperforming previous vision and language assistants, and even achieving comparable or better performance than specialist models trained to perform these specific tasks. Furthermore, TEOChat achieves impressive zero-shot performance on a change detection and change question answering dataset, outperforms GPT-4o and Gemini 1.5 Pro on multiple temporal tasks, and exhibits stronger single image capabilities than a comparable single EO image instruction-following model. We publicly release our data, models, and code: https://github.com/ermongroup/TEOChat.
@@ 1 Introduction
Many earth observation (EO) tasks require the ability to reason over time. For example, change detection is a widely studied task where the goal is to identify salient changes in a region using multiple EO images capturing the region at different times (Chughtai et al., 2021; Bai et al., 2023; Cheng et al., 2023). Automated methods to effectively perform change detection can significantly aid humanitarian and sustainability efforts, including planning disaster relief (Rahnemoonfar et al., 2021; Gupta et al., 2019; Ban et al., 2020), tracking urban changes (Tamilenthi & Baskaran, 2013; Fyleris et al., 2022; Verma et al., 2021), and monitoring deforestation (Shermeyer & Haack, 2015; De Bem et al., 2020; Decuyper et al., 2022), crops, (Gim et al., 2020; Kaur et al., 2023; Wang et al., 2024a), and ecological conditions (Willis, 2015; Xu et al., 2019). Previous methods to automatically detect change in EO imagery have been specialist models, constraining their use to a single task or small set of tasks that they were explicitly trained to perform (Bai et al., 2023; Cheng et al., 2023).
Advancements in the modeling of multimodal data have enabled generalist vision-language models (VLMs) that can perform a variety of natural image interpretation tasks specified flexibly through natural language (Achiam et al., 2023; Team et al., 2023; Liu et al., 2023). These models have recently gained capabilities to reason over temporal sequences of natural images (video) (Maaz et al., 2023; Lin et al., 2023; Jin et al., 2023; Li et al., 2023; Zhang et al., 2023b; Liu et al., 2024b) and single EO images (Hu et al., 2023; Kuckreja et al., 2023; Zhan et al., 2024; Muhtar et al., 2024; Pang et al., 2024). However, no prior VLMs can model temporal EO data (left of Figure 1), notably including change detection tasks. We investigate the performance of Video-LLaVA (Lin et al., 2023), a strong natural image pre-trained VLM that can receive images and videos as input, and GeoChat (Kuckreja et al., 2023), a strong VLM fine-tuned on single EO image tasks (right of Figure 1). We find that Video-LLaVA generates inaccurate information, likely because it has primarily been trained on natural images and videos, whereas GeoChat can only input single images and cannot process information across time.
@@
Figure 1: Left: Capabilities of VLMs. TEOChat is the first VLM to model temporal earth observation (EO) data. Right: Example outputs of previous VLMs. We compare a temporal VLM (Video-LLaVA (Lin et al., 2023)) and an EO VLM (GeoChat (Kuckreja et al., 2023)) with TEOChat.
To address these limitations, we introduce TEOChat, the first large language and vision assistant that can handle temporal EO imagery. Specifically, our contributions are:
@@
    We introduce the first temporal EO dataset for multimodal instruction-tuning called TEOChatlas which we use to train TEOChat. TEOChatlas is composed of 554,071 examples spanning dozens of temporal instruction-following tasks.
@@
    We demonstrate that (i) TEOChat achieves impressive zero-shot performance on an EO change detection and a change QA dataset, (ii) TEOChat outperforms two strong proprietary foundation models for modeling sequences of images (GPT-4o and Gemini-1.5 Pro), and (iii) TEOChat possesses strong single image capabilities, outperforming GeoChat on multiple zero-shot scene classification and visual question answering tasks.
@@ 2 Developing a VLM for Temporal EO Data
We aim to develop a VLM that can input a natural language instruction (prompt) and a temporal sequence of EO images, and output a natural language response. For example, given a sequence of satellite images capturing an urban area before and after a hurricane, the model should be able to respond to questions about damaged buildings or flooded regions. The vast majority of VLMs that can perform this task on natural images follow a LLaVA-like architecture (Liu et al., 2024a) consisting of (i) a large language model (LLM; commonly LLaMA 2 (Touvron et al., 2023) initialized with Vicuna weights (Chiang et al., 2023)) (ii) a vision encoder (commonly a ViT (Dosovitskiy et al., 2020) pre-trained with CLIP (Radford et al., 2021)) to generate the visual representations, and (iii) a vision-language connector (commonly an MLP (Liu et al., 2023)) to align the visual representations with the natural language token embeddings input to the LLM.
Training the VLM requires aligning the language representations with the visual concepts in EO images, but these concepts are not well-represented in natural image data. EO imagery is taken from a birds-eye view and captures land use and land cover changes like flooding or road construction. To align the the EO imagery with natural language and allow the model to provide responses to instructions about the EO data, we require a dataset of triplets consisting of sample instructions, EO images, and ground truth responses, commonly referred to as an instruction-following dataset. The responses are used to supervise the VLM using a next-token cross-entropy loss on the response tokens (visual instruction tuning (Liu et al., 2024a)). We note that existing instruction-following EO datasets are insufficient for training a temporal EO VLM, as the tasks do not require temporal capabilities and therefore do not promote their development. We confirm this empirically by showing that GeoChat substantially underperforms our approach across all temporal reasoning tasks we test.
@@ 3 TEOChatlas: A Temporal EO Instruction-Following Dataset
The main obstacle to developing a temporal EO VLM is the absence of a suitable instruction-following dataset. To overcome this, we curate an instruction-following dataset of temporal EO imagery called TEOChatlas. A few key desired capabilities of the VLM induce important dataset design decisions. First, the model should be able to follow instructions about a wide variety of tasks requiring spatial and temporal reasoning capabilities, so we construct a diverse composition of instruction-following tasks (Section 3.1). Second, the model should be able to reason over variable-length temporal sequences of EO data from different sensors, so we ensure that many sensors and sequence lengths are represented (Section 3.2). Third, the model should be able to reference specific images in its input and output and support user-friendly conversation, so we design the dataset prompts to support these features (Section 3.3).
@@ 3.1 TEOChatlas: Task Composition
We curate various instruction-following tasks in TEOChatlas in order to develop a range of capabilities in the model. Importantly, because our goal is to add new temporal reasoning capabilities without compromising single image reasoning abilities, we include both single and temporal instruction-following tasks in the dataset, described in detail below.
@@ 3.1.1 Single Image Tasks
To support the development of single image capabilities like object recognition and spatial reasoning, we include many single image instruction-following tasks in TEOChatlas. Specifically, we use the GeoChat dataset (Kuckreja et al., 2023), a composition of single image instruction-following tasks including scene classification, visual question answering, referring expression, region captioning, detailed description, grounded description, and multi-turn conversation. The GeoChat dataset was used to train the GeoChat model, which exhibits strong performance on these single image tasks. Appendix A has more detail about the GeoChat dataset and instruction-following tasks.
@@ 3.1.2 Temporal Tasks
We construct several temporal instruction-following tasks using four common temporal EO datasets. We group the tasks into seven categories spanning two common real-world applications, namely disaster response and urban development monitoring. Figure 2 shows examples of the temporal tasks and Table 8 in the Appendix provides the full set of tasks and associated prompts. TEOChatlas is the first instruction-following EO dataset to have temporal tasks (Table 6).
@@
Change Detection (CD)  We curate multiple instruction-following tasks for change detection, a widely studied earth vision task where the goal is to identify changes between images of the same region over time. We include two canonical change detection tasks in the dataset, namely building damage assessment and building change detection.
@@
Region-based Change Question Answering (RQA)  We also create question answering tasks based on region prompt inputs to improve reasoning about spatial inputs. We generate the same tasks from xBD and S2Looking as used in change question answering, but constrain the question to a bounding box input representing a specific area within the image.
We also create tasks using QFabric (Verma et al., 2021), a dataset for multi-task urban change detection consisting of sequences of five images, where every sequence is labeled with several polygons representing urban change. Each polygon has a sequence-level label indicating the broad type of change (change type classification) and an image-level label indicating the specific temporal change that occurred between adjacent timesteps (change status classification). We construct instruction-following tasks for these two tasks by mirroring the building damage classification task, instructing the model to classify the change type or change status of an input bounding box into the label categories. Following the other classification tasks, we include the label categories in the prompt.
Temporal Referring Expression (TRE)  To enable the model to temporally localize change, we include multiple tasks instructing the model to identify the specific image(s) in which a given change occurred. We use the change type labels from QFabric which indicate the type of urban development that has occurred in each image in the temporal sequence. Because these labels are tied to specific regions of the images, we include the associated bounding box in the prompt and task the model with identifying the image(s) in the sequence where the change occurred. To perform this task, the model must be able to model spatial inputs and generate temporal outputs.
@@ 3.2 TEOChatlas: EO Image Sequences
To allow the model to reason about varying input sequence lengths and sensor types, we ensure that TEOChatlas includes a diverse representation of image sequence sizes and image providers. We note that TEOChatlas also has considerable geographic diversity (Figure 5). We provide more detail about how we process the images of each dataset in Appendix A.
For sequence lengths, the temporal EO datasets we include consist of bitemporal (xBD and S2Looking), pentatemporal (QFabric), and multitemporal (variable length) sequences (fMoW). Due to memory constraints, we limit the sequences to a maximum of 8 images, and randomly sample 8 images without replacement from longer sequences. To increase the representation of other sequence lengths in the dataset, we also randomly sample sequences of shorter lengths from QFabric.
For sensor types, the EO datasets in TEOChatlas span 8 sensors (Table 7). To increase the representation of Sentinel-2, a publicly available sensor with global coverage that is commonly used for EO tasks, we include Sentinel-2 image sequences from fMoW-Sentinel (Cong et al., 2022).
@@ 3.3 TEOChatlas: Prompt Design
We design the prompts in TEOChatlas to enable the model to reference images and converse in a user-friendly manner. The temporal referring expression and region-based temporal question answering tasks require the model to input and output references to specific images in the sequence. To enable this, we interleave image identifiers among the visual tokens in the sequence input to the LLM (see Figure 3). This differs from prior methods for modeling temporal sequences of natural images, which simply concatenate the visual representations (Maaz et al., 2023; Lin et al., 2023; Zhang et al., 2023b; Jin et al., 2023; Li et al., 2023; Touvron et al., 2023).
The original GeoChat dataset depends on task tokens to help the model differentiate between tasks, for example to determine whether to include bounding boxes in its response. We remove this requirement, and replace the tokens with short natural language specifications of the task. Examples of this are provided in Table 8 in the Appendix.
@@ 4.1 TEOChat Architecture
We adopt a LLaVA-1.5-like architecture (Liu et al., 2023) consisting of (i) a temporally-shared image encoder (CLIP ViT-L/14 (Radford et al., 2021)) to obtain representations of each image in the sequence, (ii) a 2-layer MLP to project the visual representations to the input of the LLM, and (iii) an LLM decoder (Llama 2 (Touvron et al., 2023)) which inputs the instruction and temporal sequence of projected visual representations to generate the response (Figure 3). A recent extension of LLaVA-1.5 to handle videos called Video-LLaVA (Lin et al., 2023) uses joint natural image-video instruction-tuning by using the frozen CLIP ViT image encoder for images and a frozen LanguageBind (CLIP-aligned) video encoder for videos (Zhu et al., 2023), while keeping the rest of the architecture the same as LLaVA-1.5. We test whether these projector and LLM weights learned on video data are better for temporal EO data by comparing Video-LLaVA initialization to LLaVA initialization, as well as freezing vs. fine-tuning the projector, in our experimental ablations.
Another key architectural decision is whether to use the image encoder or video encoder from Video-LLaVA (Lin et al., 2023) to generate visual representations for the sequence. We choose to use the image encoder for two reasons: (1) prior change detection approaches have demonstrated that Siamese encoders with weights shared across time (temporal-agnostic) are highly effective for identifying changes in sequences of EO data (Zheng et al., 2021a; b; 2022), and (2) the video encoder was designed to input fixed sequences of 8 images. While it is possible to remove this second constraint, we observed performance degradation when using shorter sequences in preliminary experiments, likely because the video encoder was trained strictly on 8 image sequences (Zhu et al., 2023). Replicating images in the sequence may address this but makes training and inference much slower. Altogether, our vision encoder design is fast, memory efficient, and achieves strong performance.

### v2

@@ Abstract
Large vision and language assistants have enabled new capabilities for interpreting natural images. These approaches have recently been adapted to earth observation data, but they are only able to handle single image inputs, limiting their use for many real-world tasks. In this work, we develop a new vision and language assistant called TEOChat that can engage in conversations about temporal sequences of earth observation data. To train TEOChat, we curate an instruction-following dataset composed of many single image and temporal tasks including building change and damage assessment, semantic change detection, and temporal scene classification. We show that TEOChat can perform a wide variety of spatial and temporal reasoning tasks, substantially outperforming previous vision and language assistants, and even achieving comparable or better performance than several specialist models trained to perform specific tasks. Furthermore, TEOChat achieves impressive zero-shot performance on a change detection and change question answering dataset, outperforms GPT-4o and Gemini 1.5 Pro on multiple temporal tasks, and exhibits stronger single image capabilities than a comparable single image instruction-following model on scene classification, visual question answering, and captioning. We publicly release our data, model, and code at https://github.com/ermongroup/TEOChat.
@@ 1 Introduction
Many earth observation (EO) tasks require reasoning over time. For example, change detection is a widely studied task with the goal of identifying salient changes in a region using EO images at different times (Chughtai et al., 2021). Automated methods to perform change detection can greatly aid humanitarian and sustainability efforts, including planning disaster relief (Khan et al., 2023), tracking urban change (Reba & Seto, 2020), and monitoring deforestation (Mitchell et al., 2017), crops (Karmakar et al., 2024), and ecological conditions (Besson et al., 2022). Prior automated methods to detect change in EO imagery have been specialist models, constraining their use to single tasks they were explicitly trained to perform (Bai et al., 2023; Cheng et al., 2024).
Advancements in the modeling of multimodal data have enabled generalist vision-language models (VLMs) that can perform a variety of natural image interpretation tasks specified through natural language (Liu et al., 2023a). These models combine the capabilities of large language models and vision encoders, aligning visual and textual representations to enable natural, conversational interfaces for diverse tasks. This flexible interface presents a major opportunity for developing multimodal conversational agents for a variety of real-world applications, such as emergency response, where rapid and adaptable information processing is crucial.
New VLMs have gained capabilities to model temporal sequences of natural images (video) (Tang et al., 2023) and single EO images (Li et al., 2024c). However, no prior VLMs can model temporal EO data (left of Figure 1). We investigate the performance of Video-LLaVA (Lin et al., 2023), a strong natural image VLM that can receive images and videos as input, and GeoChat (Kuckreja et al., 2023), a strong VLM fine-tuned on single EO image tasks (right of Figure 1). We find Video-LLaVA generates inaccurate information, likely due to being trained on natural images and videos, whereas GeoChat only inputs single images and cannot process information across time.
We address these limitations by introducing TEOChat, the first VLM for temporal sequences of EO images. TEOChat adapts the Video-LLaVA architecture, designed originally for natural images and videos, to single and temporal sequences of EO images. To train TEOChat, we curate TEOChatlas, the first instruction-tuning dataset with instruction-following examples for temporal EO data. We construct a variety of tasks which require spatial and temporal reasoning capabilities using four EO datasets, namely fMoW (Christie et al., 2018), xBD (Gupta et al., 2019), S2Looking (Shen et al., 2021), and QFabric (Verma et al., 2021). TEOChatlas is large and diverse in task composition, sensors, and geography, making it a strong dataset for training generalizable EO VLMs.
@@
Figure 1: Left: VLM Capabilities. TEOChat is the first VLM for temporal EO data. Right: Example outputs of prior VLMs. We compare to a temporal VLM (Video-LLaVA) and a single EO image VLM (GeoChat).
In summary, our contributions are:
@@
    We introduce the first temporal EO dataset for multimodal instruction-tuning called TEOChatlas to train TEOChat. TEOChatlas contains 245,210 temporal EO examples spanning dozens of temporal instruction-following tasks as well as 308,861 single EO examples from GeoChat_Instruct, totalling 554,071 examples and 1,244,393 images.
@@
    We conduct a thorough set of experimental ablations to motivate TEOChat’s design. We find the key design decisions include Video-LLaVA initialization, vision-language connector fine-tuning, and the inclusion of image references between visual tokens in the prompt.
  4. 4.
    We demonstrate that (i) TEOChat achieves impressive zero-shot performance on an EO change detection and a change QA dataset, (ii) TEOChat outperforms two strong proprietary foundation models for modeling sequences of images (GPT-4o and Gemini-1.5 Pro), and (iii) TEOChat possesses strong single image capabilities, outperforming GeoChat on multiple zero-shot scene classification and visual question answering (VQA) tasks.
@@ 2 Developing a VLM for Temporal EO Data
We aim to develop a VLM that can input a natural language instruction and a temporal sequence of EO images, and output a natural language response. For example, given satellite images capturing an urban area before and after a hurricane, the model should respond to questions about damaged buildings or flooded regions. The vast majority of natural image VLMs for this task use a LLaVA-like architecture (Liu et al., 2024a) consisting of (i) a large language model (LLM; commonly LLaMA 2 (Touvron et al., 2023) initialized with Vicuna weights (Chiang et al., 2023)) (ii) a vision encoder (commonly a ViT (Dosovitskiy et al., 2020) pre-trained with CLIP (Radford et al., 2021)) to generate visual representations, and (iii) a vision-language connector (commonly an MLP (Liu et al., 2023a)) to align the visual representations with the language embeddings input to the LLM.
Training the VLM requires aligning language representations with visual concepts in EO images, but these concepts are not well-represented in natural image data. EO imagery is taken from a birds-eye view and captures land use and land cover changes like flooding or road construction. To align EO data with language and allow the model to respond to instructions about it, we require a dataset of instruction, EO image, and ground truth response triplets (an instruction-following dataset). The responses are used to supervise the VLM using a next-token cross-entropy loss on the response tokens (visual instruction tuning (Liu et al., 2024a)). We note that existing instruction-following EO datasets are insufficient for training a temporal EO VLM, as the tasks do not require temporal capabilities and therefore do not promote their development. We confirm this empirically by showing that GeoChat substantially underperforms our approach across all temporal reasoning tasks we test.
@@ 3 TEOChatlas: A Temporal EO Instruction-Following Dataset
The main obstacle to developing a temporal EO VLM is the absence of a suitable instruction-following dataset. To overcome this, we curate an instruction-following dataset of temporal EO imagery called TEOChatlas. A few key desired capabilities of the VLM induce important dataset design decisions. First, the model should follow instructions about a wide variety of tasks requiring spatial and temporal reasoning capabilities, so we construct a diverse set of instruction-following tasks (Section 3.1). Second, the model must reason over variable-length temporal sequences of EO data from different sensors, so we ensure many sensors and sequence lengths are represented (Section 3.2). Third, the model should be able to reference specific images in its input and output, and support user-friendly conversation, so we design the prompts to support these features (Section 3.3).
@@ 3.1 TEOChatlas: Task Composition
We curate instruction-following tasks in TEOChatlas to develop a range of model capabilities. Importantly, because our goal is to add temporal capabilities without compromising single image capabilities, we include both single image and temporal tasks in the dataset, described below.
@@ 3.1.1 Single Image Tasks
To support the development of single image capabilities like object recognition and spatial reasoning, we include many single image instruction-following tasks in TEOChatlas. Specifically, we use GeoChat_Instruct (Kuckreja et al., 2023), a composition of single image instruction-following tasks including scene classification, visual question answering, referring expression, region captioning, detailed description, grounded description, and multi-turn conversation. GeoChat_Instruct was used to train the GeoChat model, which exhibits strong performance on these single image tasks. Appendix B has more detail about GeoChat_Instruct and instruction-following tasks.
@@ 3.1.2 Temporal Tasks
We construct temporal instruction-following tasks using four EO datasets. We group tasks into seven categories spanning two common real-world applications, namely disaster response and urban development monitoring. Figure 2 shows example of tasks and Table 8 enumerates all of the tasks and prompts. TEOChatlas is the first instruction-following EO dataset with temporal tasks (Table 6).
@@
Change Detection (CD)  We curate instruction-following tasks for change detection, a widely studied earth vision task to identify changes in images of an area over time. We include two canonical change detection tasks, namely building damage assessment and building change detection.
@@
Region-based Change Question Answering (RQA)  We create question answering tasks based on regions to improve reasoning about spatial inputs. We generate the same change question answering tasks from xBD and S2Looking, but constrain the question to a bounding box input representing a specific area within the image. We also create tasks using QFabric (Verma et al., 2021), a multi-task urban change detection dataset consisting of five image sequences each labeled with polygons representing urban change. Each polygon has a sequence-level label indicating the broad type of change (change type classification) and an image-level label indicating the specific change that occurred between adjacent timesteps (change status classification). We construct instruction-following tasks by mirroring building damage classification, instructing the model to classify the change type or change status of an input bounding box into categories included in the prompt.
Temporal Referring Expression (TRE)  To enable the model to temporally localize change, we include tasks instructing the model to identify the specific image(s) in which a given change occurred. We use the change type labels from QFabric which indicate the type of urban development that occurred in each image. Because these labels are tied to specific regions of the images, we include a bounding box in the prompt and task the model with identifying the image(s) where the change occurred. To perform this task, the model has to model spatial inputs and generate temporal outputs.
@@ 3.2 TEOChatlas: EO Image Sequences
To allow the model to reason about varying input sequence lengths and sensor types, we ensure that TEOChatlas includes a diverse set of image sequence sizes and image providers. For sequence lengths, the temporal EO datasets we include consist of bitemporal (xBD and S2Looking), pentatemporal (QFabric), and multitemporal (variable length) sequences (fMoW). Due to memory constraints, we limit the sequences to a maximum of 8 images, and randomly sample 8 images without replacement from longer sequences. To increase the representation of other sequence lengths in the dataset, we also randomly sample shorter sequences from QFabric. For sensor types, the EO datasets in TEOChatlas span 8 sensors (Table 7). To increase the representation of Sentinel-2, a publicly available sensor with global coverage that is commonly used for EO tasks, we include sequences from fMoW-Sentinel (Cong et al., 2022). More detail about how we process the images is in Appendix B, and we also note that TEOChatlas has considerable geographic diversity (Figure 5).
@@ 3.3 TEOChatlas: Prompt Design
We design the prompts to enable the model to reference images and converse in a user-friendly manner. As the TRE and RTQA tasks require the model to input and output references to specific images in the sequence, we design the prompts to facilitate this. To do so, we interleave image references between the visual tokens input to the LLM (see Figure 3). This differs from prior methods for temporal sequences of natural images which simply concatenate the representations (Maaz et al., 2023; Lin et al., 2023; Zhang et al., 2023b; Jin et al., 2023; Li et al., 2023; Touvron et al., 2023).
GeoChat_Instruct depends on task tokens to help the model differentiate between tasks, for example to determine whether to include boxes in its response. We remove this requirement, and replace the tokens with short natural language specifications of the task. Examples are provided in Table 8.
@@ 4.1 TEOChat Architecture
We adopt a LLaVA-1.5-like architecture (Liu et al., 2023a) consisting of (i) a temporally-shared image encoder (CLIP ViT-L/14 (Radford et al., 2021)) to obtain representations of each image in the sequence, (ii) a 2-layer MLP to project the visual representations to the input of the LLM, and (iii) an LLM decoder (Llama 2 (Touvron et al., 2023)) which inputs the instruction and temporal sequence of projected visual representations to generate the response (Figure 3). A recent extension of LLaVA-1.5 to handle videos called Video-LLaVA (Lin et al., 2023) uses joint natural image-video instruction-tuning by using the frozen CLIP ViT image encoder for images and a frozen LanguageBind (CLIP-aligned) video encoder for videos (Zhu et al., 2023a), while keeping the rest of the architecture the same as LLaVA-1.5. We test whether these projector and LLM weights learned on video data are better for temporal EO data by comparing Video-LLaVA initialization to LLaVA initialization, as well as freezing vs. fine-tuning the projector, in our experimental ablations.
Another key architectural decision is whether to use the image encoder or video encoder from Video-LLaVA (Lin et al., 2023) to generate visual representations for the sequence. We choose to use the image encoder for two reasons: (1) prior change detection approaches have demonstrated that Siamese encoders with weights shared across time (temporal-agnostic) are highly effective for identifying changes in sequences of EO data (Zheng et al., 2021a; c; 2022a), and (2) the video encoder was designed to input fixed sequences of 8 images. While it is possible to remove this second constraint, we observed performance degradation when using shorter sequences in preliminary experiments, likely because the video encoder was trained strictly on 8 image sequences (Zhu et al., 2023a). Replicating images in the sequence may address this but makes training and inference much slower. Altogether, our vision encoder is fast, memory efficient, and achieves strong performance.